Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day05


Crew 101 Log Book for March 16, 2011

Photos of the day





Commander's Check-In Report

Tom Caillouet Reporting

Crew Health: Mostly Good, Slowly returning to good health, all

others in fine condition.

EVAs and Times: No EVAs

Number of Reports: 5 (Commander, Engineering, Journalist, Chef, Science Report)

Field Results: N/A

Plans for Tomorrow: Continuing geology experiment, beginning areohashing activities. Possibly answering questions over email for Georgia Tech presentation of our current activities (11 am EST). Voice chat with NASA INSPIRE students at 9 pm EST


Commander's Report

Tom Caillouet Reporting

Today has been a bit slow of a day. After yesterday's issues with the ATVs, outside help was due to arrive at some point in the day. A small break in sim, but now our science collection will be far more sample sites available. Unfortunately, this lead to a cancellation of all EVAs planned for the day (more geology sample collection). This left us with plenty of time to analyze  EVA 4's samples and continue to improve the reliability of the radio system. Tomorrow, we return to our regularly scheduled sample collections and begin the areohashing scouting.

A note on crew health: All crewmembers besides me remain in good health. I am still not 100%, but the flight surgeon  has confirmed that it probably just dehydration. The transition from living in such a humid area to spending extended time in what feels like the driest place on Earth (err... Mars) has not been pleasant.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level

o V everstart

o V armor plate

o V green

Notes/Comments (include how many generator runs, approximately when,

and any times you needed to turn the generator on/off):

1 generator run


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.75

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

Honey Oil Quantity (# of Quarts in storage): Leave it to DG

ATV Oil Quantity (# of Quarts in storage): 6 quarts on the

engineering work bench

Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:


Water Status:

Outside Potable Water Tank Level (inches from bottom): 9

Trailer Potable Water Tank Level (inches from bottom): 19

Hab Potable Water Tank Level (inches from bottom): 13

Potable Water Meter Reading: 435712

Notes/Comments :


GreenHab:

Condition of plants in Tank 1: unchanged

Condition of plants in Tank 2: unchanged

Condition of plants in Tank 3: unchanged

Greenhab notes/ comments:

Internet still bad. Cannot attach pictures.


Observatory:

• Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): No

Oil Checked (yes/no): No

Tire Status: Not checked

Odometer at end of day: 123688.4

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: no

Oil Checked: no

Fuel Consumed: none

Tire Status: ok

Comments:


ATV 2 (Viking I):

Used: yes, started

Oil Checked: no

Fuel Consumed: negligible

Tire Status: ok

Comments:


ATV 3 (Viking II):

Used: yes, started

Oil Checked: no

Fuel Consumed: negligible

Tire Status: ok

Comments: battery was dead, and there was oil all over top of

engine. DG fixed when he stopped by.


ATV 4 (Spirit):

Used: yes, started

Oil Checked: no

Fuel Consumed: negligible

Tire Status: ok

ATV Comments: DG stopped by to show us how to start ATVs


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 68

Actual temperature upstairs (Farenheit): 72

Thermostat setting downstairs (Farenheit): 55

Actual temperature downstairs (Farenheit): 67

Notes/Comments:


Computers/Networking Infrastructure: working better, but still goes

out occasionally

Notes/comments on internet/computers:


General Engineering

Notes/Comments: We turned on the hot water heater because DG asked us to look for a leak when it was on. After some time, we noticed the water pump on the second floor of the hab turning on every other second. Upon investigation, we noticed hot water coming out of the hot water heater vent on the side of the hab. After a few minutes the pump turned on less and less until it stopped. I assume for some reason the hot water heater is set to empty out the side when it is full, but that’s just a guess.



Science Report (Biology)

Christina Graves and Christine Redmond Reporting

On 3-15-11, an on-foot EVA was performed by Commander Calliouet and Christina Graves.  The purpose of the EVA was to collect soil samples and geological data from three separate sites of the San Rafael Swell (the site of the MDRS) for analysis.  The San Rafael Swell is characterized by both anticline and sedimentary rocks. Geological data obtained includes rock morphology, type, and location.

Data obtained from the EVA is found below:

Site

Local Time

GPS Location Data

Similarity (rocks >6in)

Depth (in)

Description



UTM

 

Acc (m)

Alt (m)


1

13:57

12S 0518290

4250809

7

 

3

2

flat area, northeast of hab




2

14:02

12S 0518342

4250833

2

1370

13

0.5

near water runoff, hilly site




3

14:13

GPS battery death

0

4

large rock desert outcropping




At Site 1, a 6’ x 6’ area was marked at 12S 0518290  4250809 UTM; with an accuracy of 7m. The 108.33g sample collected at this site was a uniform sample, composed entirely of siltstone.  The sample was collected at a depth of 1.”  

At Site 2, a 6’ x 6’ area was marked at 12S 0518342  4250833 UTM; with an accuracy of 2m, at an elevation of 1370 m.  The 220.98g sample contained siltstone, some (carbonaceous?) black rock, as well as some shale.  Of particular interest, the location selected was near a hilly site, apparently a water-run off.  The site was chosen because of the probability of bacterial variety at the site.  The site also serves as a potential model of a “primordial pool,” thought to be necessary for life in prebiotic situations.

At Site 3, a 6’ x 6’ area was marked at an area approximately 200 feet E from site 2, at a large rock desert outcropping, on top of a small hill.  A 464.64g sample was collected at a depth of 4”. Small multi-colored cobblestones were found at this site in addition to a large amount of sand and siltstone found at the 2 previous sites.  

In conclusion, similar types of rock were found at the 3 sites.  This is in part due to the fact that the EVA was performed on foot, and therefore the sites were limited by range. 

Biological soil analysis (E.Z.N.A. Soil Analysis [Omega]) will be performed at a later time to test for the presence of bacteria, and other carbonaceous compounds.


Chef's Report

Christine Redmond Reporting

1a) Today was a pre-prepared day

1b) Was today a special day celebrated at the main meal with special

food or activity?

no

2) List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Mountain Chili

Broccoli and Cauliflower

3) List any main meal foods not finished by the crew (leftovers stored

for later use or discarded)

none

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

none

5) Lessons learned (if any)

Use three cups of water instead of six.

6) Comments/questions for Kim and Jean, the food study investigators


Journalist Report

Christine Redmond Reporting!

Today there was a wind storm on Mars! The crew remained grounded at base due to the ATVs being down and the windy conditions. However, crew 101 made a productive “in hab” day. This morning we prepared for our outreach chat with the NASA INSPIRE students. We did a test chat and measured the bandwidth usage after the chat. Crew 101 has been especially careful with our internet usage since the outage on Monday. We are all very excited for the chat tomorrow night!

Jon had a busy day tweaking the radio equipment. Christina and I classified the geological samples from yesterday’s EVA. Tommy had a busy afternoon updating the website for our friends back on earth to enjoy. This afternoon Don made a visit and helped us get the ATVs back up and running. We are all excited that they are back in working order and we are ready to head out on some exciting EVAs!